OpenStudy (anonymous):

find log[2] 90, if log[2] 5=2.32

OpenStudy (asnaseer):

the closest I can get is:\[\log_2(90)=\log_2(2*3^2*5)=\log_2(2)+\log_2(3^2)+\log_2(5)\]\[=1+2\log_2(3)+2.32=3.32+2\log_2(3)\]